. The Pointer Appliance Company is investigating the addition of a new and improved pulsating blender to its line of consumer appliances. The product chops, grinds, grates and blends smoothies twice as fast as all other blenders on the market. It has a 10-year warranty on the motor and offers a replacement 5-cup glass blender for $10.99. The blender is going to require additional expenditures for the production line on the part of the Pointer Appliance Company. Management wants to determine what to price the blender at. The total fixed cost associated with manufacturing the blenders is $525,000. The variable cost is $17.99 per unit. Given the costs for adding the blender, management is considering selling the blender at one of these prices: $25.99, $34.99 and $49.99. Calculate the breakeven point (BEP) in units at each of the selling prices: $25.99, $34.99 and $49.99 and show the calculations and the BEP in units at each price.

BEP units at $25.99

BEP units at $34.99.

BEP units at $49.99=

Through your research you know that this is a tough competitive market with many competitors. Most competitors have between 2-8% of the market and the market appears to be growing at a 5% rate per year. The total market for blenders of this type was 1,500,000 units last year. Given that the market leader is selling its blender for $39.99, which price would you recommend and why?
Answer #1

Answer:

Total Fixed cost = $ 525,000

Variable Cost = $ 17.99 per unit

Let X be the break-even units

Calculation of Break-even units at a price level of $ 25.99

$ 25.99 * X = $ 17.99 *X + $ 525,000

X *($25.99-$17.99) = $ 525,000

X * $ 8 = $ 525,000

X = $ 525,000/$ 8 = 65,625

Break-even level = 65,625 Units

Calculation of Break-even units at a price level of $ 34.99

$ 34.99 * X = $ 17.99 *X + $ 525,000

X *($34.99-$17.99) = $ 525,000

X * $ 17 = $ 525,000

X = $ 525,000/$ 17 = 30,882.35   or 30,883 (rounded off)

Break-even level = 30,883 Units

Calculation of Break-even units at a price level of $ 49.99

$ 49.99 * X = $ 17.99 *X + $ 525,000

X *($49.99-$17.99) = $ 525,000

X * $ 32 = $ 525,000

X = $ 525,000/$ 32 = 16,406.25 or 16,407 (rounded off)

Break-even level = 16,407 Units

Total Market Size last year = 1,500,000

Annual growth rate of market = 5%

Expected market size next year = 1,500,000 * 1.05 = 1,575,000

Assuming the company is targettting for a 6% (average of 2-8%) market share, the company needs to sell 1,575,000 * 6% = 94,500 units next year.

Calculation of profitability at various price levels for 94500 units is as follows

Calculation of profit

Sales Price $ 25.99

$ 34.99

$ 39.99 (price of market leader)

Sales

$2,456,055

$ 3,306,555

$ 3,779,055

Variable cost

$1,700,055

$ 1,700,055

$ 1,700,055

Fixed Costs

$ 525,000

$ 525,000

$ 525,000

Net Margin

$ 231,000

$ 1,081,500

$ 1,554,000

  

As the company is trying to capture market share, they cannot sell at $ 49.99 which is higher than the price of market leader or at the price of market leader which is $ 39.99. Hence they need to choose between $ 25.99 and $ 34.99. As the profitability is better at $ 34.99, they may consider selling at this price.